# Modeling Domain and Feedback Transitions for Cross-Domain Sequential Recommendation

### v1

@@
Abstract
Nowadays, many recommender systems encompass various domains to cater to users’ diverse needs, leading to user behaviors transitioning across different domains. In fact, user behaviors across different domains reveal changes in preference toward recommended items. For instance, a shift from negative feedback to positive feedback indicates improved user satisfaction. However, existing cross-domain sequential recommendation methods typically model user interests by focusing solely on information about domain transitions, often overlooking the valuable insights provided by users’ feedback transitions. In this paper, we propose $\text{Transition}^{2}$, a novel method to model transitions across both domains and types of user feedback. Specifically, $\text{Transition}^{2}$ introduces a transition-aware graph encoder based on user history, assigning different weights to edges according to the feedback type. This enables the graph encoder to extract historical embeddings that capture the transition information between different domains and feedback types. Subsequently, we encode the user history using a cross-transition multi-head self-attention, incorporating various masks to distinguish different types of transitions. Finally, we integrate these modules to make predictions across different domains. Experimental results on two public datasets demonstrate the effectiveness of $\text{Transition}^{2}$.
Introduction
@@
Figure 1: A toy illustration and statistics of domain and feedback transitions in cross-domain sequential recommendation. Type 1 transition: when transitioning across domains, the user feedback changed from positive to negative; Type 2 transition: when the user browses the new domain, her feedback improved from negative to positive.
Traditional sequential recommendation models (Zhang et al. 2024; Kang and McAuley 2018) a user’s history within a single domain, such as music (Zhang et al. 2022) or videos (Zhao et al. 2024). However, modern recommender systems often encompass multiple domains, such as books (Anwar, Siddiqui, and Saquib Sohail 2019) and movies (Goyani and Chaurasiya 2020), leading users to transition between different domains. In this transition process, not only is the information about the domain transitions valuable (referred to domain transition), but the changes in user preference toward recommended items during domain transitions are also crucial (termed feedback transition).
We classify users’ feedback transitions during their domain transitions into two types: (1) Type 1 transition: when transitioning across domains, the user feedback changed from positive to negative; (2) Type 2 transition: when browsing the new domain, the user’s feedback improves from negative to positive. For example, as shown in Figure 1(a), for the transition of “Type 1”, a user may search for related books after watching a movie she likes. However, since the movie is adapted from a book, directly recommending the original book might be disliked because she has already watched the movie. In this case, the ideal recommendation would be the sequel to that book. For a “Type 2” transition, the user is dissatisfied with several consecutively recommended books and turns to watch movies instead. This also suggests that when a user is consistently dissatisfied with items in one domain, it may be beneficial to recommend items from another domain.
@@
Existing work typically focuses on modeling users’ domain transition behaviors, to improve the performance of cross-domain sequential recommendations. A pioneering work transfers single-domain representations learned from a single domain to other domains using a gated transfer module (Ma et al. 2019). Recent work models single-domain and cross-domain sequences separately and generates representations through a self-attention mechanism (Cao et al. 2022; Ma et al. 2024). Despite the effectiveness of existing works, they overlooked the impact of users’ feedback transitions during cross-domain behaviors, and often only focused on the positive feedback from users.
To model both the domain transitions and feedback transitions in users’ cross-domain behaviors, this paper proposes an approach named $\text{Transition}^{2}$, which models domain and feedback transitions for the cross-domain sequential recommendation. Firstly, we encode users’ mixed histories across different domains using the transition-aware graph encoder. Specifically, when constructing the graph, we connect consecutive items in the history and assign different weights to the edges based on the feedback between the two items. This allows us to capture transitions between different domains and different types of feedback. After encoding with the graph encoder, we obtain item embeddings that capture transitions between different domains and feedback types. We then input these history embeddings into a cross-transition multi-head self-attention to further model the transition information in the user’s history. To model the transitions between different domains and feedback types, we introduce different masks into the heads of the transformer’s multi-head self-attention. This allows each head to capture different types of transitions. We then fuse the outputs from different heads to obtain the final history representations which include various transition information. Finally, we use this representation for prediction tasks in different domain recommendations.
The major contributions of the paper are as follows:
@@
  • We proposed $\text{Transition}^{2}$, which models domain and feedback transitions for cross-domain sequential recommendations using a GNN and a transformer equipped with different masks.
@@
Cross-Domain Sequential Recommendation (CDSR) addresses the challenge of recommending items across diverse domains by leveraging various methodological approaches. Perera and Zimmerman pioneered using timestamp information to divide user interactions into temporal itemsets, enhancing the modeling of users’ dynamic interests across domains (Perera and Zimmermann 2020a, b). Zhang et al. further advanced this by proposing the CGN model, utilizing dual generators to map itemsets across domains concurrently (Zhang et al. 2020). Another strand of research focuses on domain-specific sequential modeling, exemplified by $\pi$-Net (Ma et al. 2019) and DASL (Li et al. 2021), which employ GRUs and attention mechanisms for knowledge transfer. Recent innovations like DA-GCN (Chen et al. 2019) and MIFN (Ma et al. 2022) introduce graph-based approaches to link domain-specific item sequences, while industry applications such as SEMI (Lei et al. 2021) and RecGURU (Li et al. 2022) apply adversarial learning and multi-modal data fusion for cross-domain short-video recommendations. Novel hybrid models like C2DSR (Cao et al. 2022) integrate graphical and attentional mechanisms, employing contrastive objectives to enhance both intra-domain and cross-domain user representations. DREAM (Ye, Li, and Yao 2023) focuses on modeling decoupled representations for both single- and cross-domain. In summary, current CDSR methods fail to effectively extract and integrate cross-domain transition information, disregarding user negative feedback and thus neglecting the transition of information between positive and negative feedback. These challenges are crucial in cross-domain recommendation scenarios for predicting future user interactions.
Figure 2: The overall framework of $\text{Transition}^{2}$ consists of four modules: (1) Embedding Layer for embedding initialization, (2) Transition-Aware Graph Encoder for graph representation of user sequence considering transition, (3) Cross-Transition Multi-Head Self-Attention for capturing transition information with various masks, (4) Prediction Module for optimization.
Task Formulation
@@
In this paper, we discuss a comprehensive CDSR scenario where each user history involves two domains, namely domain $A$ and $B$. Each instance $S_{A},S_{B},S_{C}$ corresponds to a specific user. For any given instance, $S_{A}=[(a_{1},y_{1}),\ldots,(a_{|S_{A}|},y_{|S_{A}|})]$ and $S_{B}=[(b_{1},y^{\prime}_{1}),\ldots,(b_{|S_{B}|},y^{\prime}_{|S_{B}|})]$ represent the single-domain user histories, where $y_{i}$ and $y^{\prime}_{i}$ indicate the feedback associated with $a_{i}$ and $b_{i}$, respectively, taking values of $+1$ (positive feedback) or $-1$ (negative feedback). The sequence $S_{C}=[(b_{1},y^{\prime}_{1}),(a_{1},y_{1}),\ldots,(b_{|S_{B}|},y^{\prime}_{|S_{B}|}),\ldots,(a_{|S_{A}|},y_{|S_{A}|})]$ represents the cross-domain user history, formed by merging $S_{A}$ and $S_{B}$ in chronological order, where each $a\in A$ and $b\in B$ are the interacted items, and ||$\cdot$|| ⋅ | denotes the total number of items. Note that $A$ and $B$ denote the entire item sets in domain $A$ and domain $B$, respectively. Given the observed user history $(S_{A},S_{B},S_{C})_{u}$, the goal of CDSR is to predict the next item:
@@
Transition2: The Proposed Algorithm
@@
Overview of Transition2
The overall framework of our model Transition2 is illustrated in Figure 2. To effectively capture domain-transition and feedback-transition information: (1) $\text{Transition}^{2}$ first initializes the embeddings for two single-domain sequences and the cross-domain sequence with three embedding layers. (2) $\text{Transition}^{2}$ then constructs transition-aware graph encoder based on users’ cross-domain and single-domain sequences. Unlike previous methods that only utilize positive feedback for graph construction, we incorporate negative feedback and adjust the weight matrix of the graph neural network accordingly. (3) After obtaining the representation of the user’s historical sequences through the graph encoder, we design a cross-transition multi-head self-attention mechanism. This mechanism calculates attention between and within different behaviors of the user’s sequence to capture transition information. (4) The predicted scores are then calculated based on the final sequence representations, optimizing both single-domain and cross-domain losses.
Transition-Aware Graph Encoder
Inspired by the advantages of GNNs in handling sequential recommendations (Wu et al. 2019; Xu et al. 2019), we leverage GNNs to transform users’ sequential behaviors into graph representations. Additionally, to effectively capture domain transitions and feedback transitions, we have introduced transition-aware graph encoder.
Embedding Initialization
For the sequential recommendation settings, we construct single-domain graphs $G_{A}$ and $G_{B}$, as well as a cross-domain graph $G_{C}$, based on all users’ historical sequences, where only item nodes are included. Consequently, we initialize the item embeddings (Qu et al. 2023, 2024) in the three graphs as $\mathbf{E}^{(0)}_{A}=[\mathbf{e}^{(0)}_{A1};\mathbf{e}^{(0)}_{A2};\cdots]\in\mathbb{R}^{{|A|}\times d}$, $\mathbf{E}^{(0)}_{B}=[\mathbf{e}^{(0)}_{B1};\mathbf{e}^{(0)}_{B2};\cdots]\in\mathbb{R}^{{|B|}\times d}$, $\mathbf{E}^{(0)}_{C}=[\mathbf{e}^{(0)}_{C1};\mathbf{e}^{(0)}_{C2};\cdots]\in\mathbb{R}^{{(|A|+|B|)}\times d}$, respectively.
Transition-Aware Graph Construction
@@
$\textbf{T}_{ij}=\begin{cases}1&\text{if }y_{i}=y_{j},\\ -1&\text{if }y_{i}\neq y_{j},\\ 0&\text{if }i\text{ and }j\text{ are not adjacent}.\end{cases}$ (1)
@@
$\mathbf{W}=\left(\begin{array}[]{cc}\mathbf{0}&\mathbf{T}\\ \mathbf{T}^{\top}&\mathbf{0}\\ \end{array}\right).$ (2)
@@
$\mathbf{\widehat{W}}=\mathbf{D}^{-\frac{1}{2}}\mathbf{W}\mathbf{D}^{-\frac{1}{2}}.$ (3)
@@
$\mathbf{e}_{i}^{(k+1)}=\sum_{j\in\mathcal{N}_{i}}\frac{\textbf{T}_{ij}}{\sqrt{\left|\mathcal{N}_{i}\right|}\sqrt{\left|\mathcal{N}_{j}\right|}}\mathbf{e}_{i}^{(k)},$ (4)
@@
$\left\{\begin{array}[]{l}\mathbf{E}^{(k+1)}_{A}=\mathbf{\widehat{W}}_{A}\mathbf{E}^{(k)}_{A},\\ \mathbf{E}^{(k+1)}_{B}=\mathbf{\widehat{W}}_{B}\mathbf{E}^{(k)}_{B},\\ \mathbf{E}^{(k+1)}_{C}=\mathbf{\widehat{W}}_{C}\mathbf{E}^{(k)}_{C}.\\ \end{array}\right.$ (5)
@@
$\left\{\begin{array}[]{l}\mathbf{E}_{A}=\frac{1}{K}\sum_{k=0}^{K}\mathbf{E}^{(k)}_{A},\\ \mathbf{E}_{B}=\frac{1}{K}\sum_{k=0}^{K}\mathbf{E}^{(k)}_{B},\\ \mathbf{E}_{C}=\frac{1}{K}\sum_{k=0}^{K}\mathbf{E}^{(k)}_{C}.\\ \end{array}\right.$ (6)
Cross-Transition Multi-Head Self-Attention
@@
$\mathbf{M}^{f}_{ij}=\begin{cases}0&\text{ if }f_{i}=f_{j},\\ 1&\text{ if }f_{i}\neq f_{j}.\end{cases}$ (7)
$\mathbf{M}^{d}_{ij}=\begin{cases}0&\text{ if }d_{i}=d_{j},\\ 1&\text{ if }d_{i}\neq d_{j}.\end{cases}$ (8)
@@
$\left\{\begin{array}[]{l}\mathbf{M}_{1}=\mathbf{M}^{f}\&\mathbf{M}^{d},\\ \mathbf{M}_{2}=\mathbf{M}^{f}\&\widetilde{\mathbf{M}}^{d},\\ \mathbf{M}_{3}=\widetilde{\mathbf{M}}^{f}\&\mathbf{M}^{d},\\ \mathbf{M}_{4}=\widetilde{\mathbf{M}}^{f}\&\widetilde{\mathbf{M}}^{d}.\end{array}\right.$ (9)
@@
$\mathbf{E}^{\prime}_{C}=\mathrm{FFN}\left(\mathrm{MHSA}\left(\widehat{\mathbf{E}}_{C},\widehat{\mathbf{E}}_{C},\widehat{\mathbf{E}}_{C},\mathbf{M}_{C}\right)\right).$ (10)
@@
$\mathrm{MHSA}\left(\widehat{\mathbf{E}}_{C},\widehat{\mathbf{E}}_{C},\widehat{\mathbf{E}}_{C},\mathbf{M}_{C}\right)=\text{Concate}\left(\text{head}_{1},\text{head}_{2},\ldots\right)W.$ (11)
@@
$\text{head}_{i}$ $=\mathrm{Attention}\left(\widehat{\mathbf{E}}_{C},\widehat{\mathbf{E}}_{C},\widehat{\mathbf{E}}_{C},\mathbf{M}_{i\%4}\right)$ (12)
@@
$\mathbf{E}^{\prime}_{A}=\mathrm{FFN}\left(\mathrm{MHSA}\left(\widehat{\mathbf{E}}_{A},\widehat{\mathbf{E}}_{A},\widehat{\mathbf{E}}_{A},\mathbf{M}_{A}\right)\right),$ (13)
$\mathbf{E}^{\prime}_{B}=\mathrm{FFN}\left(\mathrm{MHSA}\left(\widehat{\mathbf{E}}_{B},\widehat{\mathbf{E}}_{B},\widehat{\mathbf{E}}_{B},\mathbf{M}_{B}\right)\right),$ (14)
Model Training
For the final model training loss, we designed losses $\mathcal{L}_{\text{single }}^{A}$ and $\mathcal{L}_{\text{single }}^{B}$ for single-domain recommendation tasks and losses $\mathcal{L}_{\text{cross }}^{A}$ and $\mathcal{L}_{\text{cross }}^{B}$ for cross-domain recommendation tasks. These are then combined to obtain the total loss:
$\mathcal{L}_{\text{total }}=\underbrace{\mathcal{L}_{\text{single }}^{A}+\mathcal{L}_{\text{single }}^{B}}_{\text{Single-Domain Loss }}+\underbrace{\mathcal{L}_{\text{cross }}^{A}+\mathcal{L}_{\text{cross }}^{B}}_{\text{Cross-Domain Loss }}.$ (15)
where the losses are all defined as the cross-entropy loss between the predictions and the ground truth:

### v2

@@
Abstract.
Nowadays, many recommender systems encompass various domains to cater to users’ diverse needs, leading to user behaviors transitioning across different domains. In fact, user behaviors across different domains reveal changes in preference toward recommended items. For instance, a shift from negative feedback to positive feedback indicates improved user satisfaction. However, existing cross-domain sequential recommendation methods typically model user interests by focusing solely on information about domain transitions, often overlooking the valuable insights provided by users’ feedback transitions. In this paper, we propose $\text{Transition}^{2}$, a novel method to model transitions across both domains and types of user feedback. Specifically, $\text{Transition}^{2}$  introduces a transition-aware graph encoder based on user history, assigning different weights to edges according to the feedback type. This enables the graph encoder to extract historical embeddings that capture the transition information between different domains and feedback types. Subsequently, we encode the user history using a cross-transition multi-head self-attention, incorporating various masks to distinguish different types of transitions. To further enhance representation learning, we employ contrastive losses to align transitions across domains and feedback types. Finally, we integrate these modules to make predictions across different domains. Experimental results on two public datasets demonstrate the effectiveness of $\text{Transition}^{2}$.
Cross-Domain Recommendation, Graph Neural Networks
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Recommender systems
1. Introduction
@@
Figure 1. A toy illustration and statistics of domain and feedback transitions in cross-domain sequential recommendation. Type 1 transition: when transitioning across domains, the user feedback changed from positive to negative; Type 2 transition: when the user browses the new domain, her feedback improved from negative to positive.
Traditional sequential recommendation models (Zhang et al., 2024a; Kang and McAuley, 2018; Zhang et al., 2025b, 2024b; Tang et al., 2025; Zheng et al., 2023) a user’s history within a single domain, such as music (Zhang et al., 2022) or videos (Zhao et al., 2024; Zhang et al., 2025a). However, modern recommender systems often encompass multiple domains, such as books (Anwar et al., 2019) and movies (Goyani and Chaurasiya, 2020), leading users to transition between different domains. In this transition process, not only is the information about the domain transitions valuable (referred to domain transition), but the changes in user preference toward recommended items during domain transitions are also crucial (termed feedback transition).
We classify users’ feedback transitions during their domain transitions into two types: (1) Type 1 transition: when transitioning across domains, the user feedback changed from positive to negative; (2) Type 2 transition: when browsing the new domain, the user’s feedback improves from negative to positive. For example, as shown in Figure 1(a), for the transition of “Type 1”, a user may search for related books after watching a movie she likes (positive). However, since the movie is adapted from a book, directly recommending the original book might be ignored (negative) because she has already watched the movie. In this case, the ideal recommendation would be the sequel to that book. For a “Type 2” transition, the user is dissatisfied (negative) with several consecutively recommended books and turns to watch movies (positive) instead. Thus, recommending cross-domain items may benefit users consistently dissatisfied with one domain’s items.
@@
Existing work typically focuses on modeling users’ domain transition behaviors, to improve the performance of cross-domain sequential recommendations. A pioneering work transfers single-domain representations learned from a single domain to other domains using a gated transfer module (Ma et al., 2019). Recent work models single-domain and cross-domain sequences separately and generates representations through a self-attention mechanism (Cao et al., 2022; Ma et al., 2024). Despite the effectiveness of existing works, they overlooked the impact of users’ feedback transitions during cross-domain behaviors, and often only focused on the positive feedback from users.
To model both the domain transitions and feedback transitions in users’ cross-domain behaviors, this paper proposes an approach named $\text{Transition}^{2}$, which models domain and feedback transitions for the cross-domain sequential recommendation. Firstly, we encode users’ mixed histories across different domains using the transition-aware graph encoder. Specifically, when constructing the graph, we connect consecutive items in the history and assign different weights to the edges based on the feedback between the two items. This allows us to capture transitions between different domains and different types of feedback. After encoding with the graph encoder, we obtain item embeddings that capture transitions between different domains and feedback types. We then input these history embeddings into a cross-transition multi-head self-attention to further model the transition information in the user’s history. To model the transitions between different domains and feedback types, we introduce different masks into the heads of the transformer’s multi-head self-attention. This allows each head to capture different types of transitions. We then fuse the outputs from different heads to obtain the final history representations which include various transition information. Furthermore, we design contrastive losses to align transitions across different domains and feedback types, helping the model better learn representations of different transitions. Finally, we use these representations for prediction tasks in different domain recommendations. The major contributions of the paper are as follows:
@@
  • We proposed $\text{Transition}^{2}$, which models domain and feedback transitions using a GNN, a transformer equipped with different masks and a transition alignment module.
@@ 2. Related Work
Cross-Domain Sequential Recommendation (CDSR) addresses the challenge of recommending items across diverse domains by leveraging various methodological approaches. Perera and Zimmerman pioneered using timestamp information to divide user interactions into temporal itemsets, enhancing the modeling of users’ dynamic interests across domains (Perera and Zimmermann, 2020a, b). Zhang et al. further advanced this by proposing the CGN model, utilizing dual generators to map itemsets across domains concurrently (Zhang et al., 2020). Another strand of research focuses on domain-specific sequential modeling, exemplified by $\pi$-Net (Ma et al., 2019) and DASL (Li et al., 2021), which employ GRUs and attention mechanisms for knowledge transfer. Recent innovations like DA-GCN (Chen et al., 2019) and MIFN (Ma et al., 2022) introduce graph-based approaches to link domain-specific item sequences, while industry applications such as SEMI (Lei et al., 2021) and RecGURU (Li et al., 2022) apply adversarial learning and multi-modal data fusion for cross-domain short-video recommendations. Novel hybrid models like C2DSR (Cao et al., 2022) integrate graphical and attentional mechanisms, employing contrastive objectives to enhance both intra-domain and cross-domain user representations. DREAM (Ye et al., 2023) focuses on modeling decoupled representations for both single- and cross-domain. TriCDR (Ma et al., 2024) uses triple cross-domain attention and contrastive learning to model comprehensive cross-domain correlations. CA-CDSR (Yin et al., 2024) focus on item representation alignment via sequence-aware augmentation and adaptive partial alignment. In summary, current CDSR methods fail to effectively extract and integrate cross-domain transition information, disregarding user negative feedback and thus neglecting the transition of information between positive and negative feedback. These challenges are crucial in cross-domain recommendation scenarios for predicting future user interactions.
Figure 2. The overall architecture of $\text{Transition}^{2}$  comprises five key components: (1) an embedding layer, which initializes item embeddings; (2) a transition-aware graph encoder, which constructs graph-based representations of user sequences while incorporating transition information; (3) a cross-transition multi-head self-attention mechanism, designed to capture diverse transition patterns through specialized masking strategies; (4) a transition alignment module, which employs contrastive learning to align representations across different transitions; and (5) a prediction module, responsible for generating recommendations and optimizing the overall objective.
3. Task Formulation
@@
In this paper, we discuss a comprehensive CDSR scenario where each user history involves two domains $A$ and $B$. Each instance $S_{A},S_{B},S_{C}$ corresponds to a specific user. For any given instance, $S_{A}=[(a_{1},y_{1}),\ldots,(a_{|S_{A}|},y_{|S_{A}|})]$ and $S_{B}=[(b_{1},y^{\prime}_{1}),\ldots,(b_{|S_{B}|},y^{\prime}_{|S_{B}|})]$ represent the single-domain user histories, where $y_{i}$ and $y^{\prime}_{i}$ indicate the feedback associated with $a_{i}$ and $b_{i}$, respectively, taking values of $+1$ (positive feedback) or $-1$ (negative feedback). The sequence $S_{C}=[(b_{1},y^{\prime}_{1}),(a_{1},y_{1}),\ldots,(b_{|S_{B}|},y^{\prime}_{|S_{B}|}),\ldots,(a_{|S_{A}|},y_{|S_{A}|})]$ represents the cross-domain user history, formed by merging $S_{A}$ and $S_{B}$ in chronological order, where each $a\in A$ and $b\in B$ are the interacted items, and ||$\cdot$|| ⋅ | denotes the total number of items. Note that $A$ and $B$ denote the entire item sets in domain $A$ and domain $B$, respectively. Given the observed user history $(S_{A},S_{B},S_{C})_{u}$, the goal of CDSR is to predict the next item:
@@
4. Transition2: The Proposed Algorithm
@@
4.1. Overview of Transition2
The overall framework of our model Transition2 is illustrated in Figure 2. To effectively capture domain-transition and feedback-transition information: (1) $\text{Transition}^{2}$  first initializes the embeddings for two single-domain sequences and the cross-domain sequence with three embedding layers. (2) $\text{Transition}^{2}$  then constructs transition-aware graph encoder based on users’ cross-domain and single-domain sequences. Unlike previous methods that only utilize positive feedback for graph construction, we incorporate negative feedback and adjust the weight matrix of the graph neural network accordingly. (3) After obtaining the representation of the user’s historical sequences through the graph encoder, we design a cross-transition multi-head self-attention mechanism. This mechanism calculates attention between and within different behaviors of the user’s sequence to capture transition information. (4) A transition alignment module with two contrastive losses aligns intra-domain item representations while contrasting those with different feedback types. (5) The predicted scores are then calculated based on the final sequence representations, optimizing both single-domain and cross-domain losses.
4.2. Transition-Aware Graph Encoder
Inspired by the advantages of GNNs in handling sequential recommendations (Wu et al., 2019; Xu et al., 2019), we leverage GNNs to transform users’ sequential behaviors into graph representations. Additionally, to effectively capture domain transitions and feedback transitions, we have introduced transition-aware graph encoder.
4.2.1. Embedding Initialization
For the sequential recommendation settings, we construct single-domain graphs $G_{A}$ and $G_{B}$, as well as a cross-domain graph $G_{C}$, based on all users’ historical sequences, where only item nodes are included. Consequently, we initialize the item embeddings (Qu et al., 2023, 2024) in the three graphs as $\mathbf{E}^{(0)}_{A}=[\mathbf{e}^{(0)}_{A1};\mathbf{e}^{(0)}_{A2};\cdots]\in\mathbb{R}^{{|A|}\times d}$, $\mathbf{E}^{(0)}_{B}=[\mathbf{e}^{(0)}_{B1};\mathbf{e}^{(0)}_{B2};\cdots]\in\mathbb{R}^{{|B|}\times d}$, $\mathbf{E}^{(0)}_{C}=[\mathbf{e}^{(0)}_{C1};\mathbf{e}^{(0)}_{C2};\cdots]\in\mathbb{R}^{{(|A|+|B|)}\times d}$, respectively.
4.2.2. Transition-Aware Graph Construction
@@
(1) $\textbf{T}_{ij}=\begin{cases}1&\text{if }y_{i}=y_{j},\\ -1&\text{if }y_{i}\neq y_{j},\\ 0&\text{if }i\text{ and }j\text{ are not adjacent}.\end{cases}$
@@
(2) $\mathbf{W}=\left(\begin{array}[]{cc}\mathbf{0}&\mathbf{T}\\ \mathbf{T}^{\top}&\mathbf{0}\\ \end{array}\right).$
@@
(3) $\mathbf{\widehat{W}}=\mathbf{D}^{-\frac{1}{2}}\mathbf{W}\mathbf{D}^{-\frac{1}{2}}.$
@@ 4.2.3. Propagation and Aggregation
(4) $\mathbf{e}_{i}^{(k+1)}=\sum_{j\in\mathcal{N}_{i}}\frac{\textbf{T}_{ij}}{\sqrt{\left|\mathcal{N}_{i}\right|}\sqrt{\left|\mathcal{N}_{j}\right|}}\mathbf{e}_{i}^{(k)},$
@@
(5) $\left\{\begin{array}[]{l}\mathbf{E}^{(k+1)}_{A}=\mathbf{\widehat{W}}_{A}\mathbf{E}^{(k)}_{A},\\ \mathbf{E}^{(k+1)}_{B}=\mathbf{\widehat{W}}_{B}\mathbf{E}^{(k)}_{B},\\ \mathbf{E}^{(k+1)}_{C}=\mathbf{\widehat{W}}_{C}\mathbf{E}^{(k)}_{C}.\\ \end{array}\right.$
@@
(6) $\left\{\begin{array}[]{l}\mathbf{E}_{A}=\frac{1}{K}\sum_{k=0}^{K}\mathbf{E}^{(k)}_{A},\\ \mathbf{E}_{B}=\frac{1}{K}\sum_{k=0}^{K}\mathbf{E}^{(k)}_{B},\\ \mathbf{E}_{C}=\frac{1}{K}\sum_{k=0}^{K}\mathbf{E}^{(k)}_{C}.\\ \end{array}\right.$
4.3. Cross-Transition Multi-Head Self-Attention
@@
(7) $\mathbf{M}^{f}_{ij}=\begin{cases}0&\text{ if }f_{i}=f_{j},\\ 1&\text{ if }f_{i}\neq f_{j}.\end{cases}$
(8) $\mathbf{M}^{d}_{ij}=\begin{cases}0&\text{ if }d_{i}=d_{j},\\ 1&\text{ if }d_{i}\neq d_{j}.\end{cases}$
@@
(9) $\left\{\begin{array}[]{l}\mathbf{M}_{1}=\mathbf{M}^{f}\&\mathbf{M}^{d},\\ \mathbf{M}_{2}=\mathbf{M}^{f}\&\widetilde{\mathbf{M}}^{d},\\ \mathbf{M}_{3}=\widetilde{\mathbf{M}}^{f}\&\mathbf{M}^{d},\\ \mathbf{M}_{4}=\widetilde{\mathbf{M}}^{f}\&\widetilde{\mathbf{M}}^{d}.\end{array}\right.$
@@
(10) $\mathbf{E}^{\prime}_{C}=\mathrm{FFN}\left(\mathrm{MHSA}\left(\widehat{\mathbf{E}}_{C},\widehat{\mathbf{E}}_{C},\widehat{\mathbf{E}}_{C},\mathbf{M}_{C}\right)\right).$
@@
(11) $\mathrm{MHSA}\left(\widehat{\mathbf{E}}_{C},\widehat{\mathbf{E}}_{C},\widehat{\mathbf{E}}_{C},\mathbf{M}_{C}\right)=\text{Concate}\left(\text{head}_{1},\text{head}_{2},\ldots\right)W.$
@@
(12) $\text{head}_{i}$ $=\mathrm{Attention}\left(\widehat{\mathbf{E}}_{C},\widehat{\mathbf{E}}_{C},\widehat{\mathbf{E}}_{C},\mathbf{M}_{i\%4}\right)$
@@
(13) $\mathbf{E}^{\prime}_{A}=\mathrm{FFN}\left(\mathrm{MHSA}\left(\widehat{\mathbf{E}}_{A},\widehat{\mathbf{E}}_{A},\widehat{\mathbf{E}}_{A},\mathbf{M}_{A}\right)\right),$
(14) $\mathbf{E}^{\prime}_{B}=\mathrm{FFN}\left(\mathrm{MHSA}\left(\widehat{\mathbf{E}}_{B},\widehat{\mathbf{E}}_{B},\widehat{\mathbf{E}}_{B},\mathbf{M}_{B}\right)\right).$
4.4. Transition Alignment and Contrast
For the extracted three types of representations, where $\mathbf{E}^{\prime}_{A}$ represents the characterization of domain A, $\mathbf{E}^{\prime}_{B}$ represents domain B, and $\mathbf{E}^{\prime}_{C}$ denotes the mix-domain representations, we first align the item representations within the same domain across these three sets, and then conduct contrastive modeling on representations with different feedback types to capture their nuanced differences.
4.4.1. Domain Representation Alignment
Given that we have three sets of representation information, the goal is to guide the model to pull semantically equivalent representations closer within the same domain. To achieve this, we introduce InfoNCE (Oord et al., 2018) to align the domain A information in $\mathbf{E}^{\prime}_{C}$ with $\mathbf{E}^{\prime}_{A}$, and the domain B information in $\mathbf{E}^{\prime}_{C}$ with $\mathbf{E}^{\prime}_{B}$. Specifically, for $\mathbf{E}^{\prime}_{C}$, we separate domain A and domain B historical records in each user sequence through indexing to obtain $\mathbf{E}^{\prime\prime}_{A}$ and $\mathbf{E}^{\prime\prime}_{B}$. We then apply mean-pooling along the sequence dimension to derive: $\mathbf{e}^{\prime\prime}_{A}=\text{MEAN}(\mathbf{E}^{\prime\prime}_{A})$ and $\mathbf{e}^{\prime\prime}_{B}=\text{MEAN}(\mathbf{E}^{\prime\prime}_{B})$. Similarly, we perform mean-pooling for $\mathbf{E}^{\prime}_{A}$ and $\mathbf{E}^{\prime}_{B}$ with $\mathbf{e}^{\prime}_{A}=\text{MEAN}(\mathbf{E}^{\prime}_{A})$ and $\mathbf{e}^{\prime}_{B}=\text{MEAN}(\mathbf{E}^{\prime}_{B})$. The alignment loss is defined using in-batch InfoNCE loss: